OpenStudy (littlebirdy):

1.The line y − 6 = 2(x + 4) passes through which one of the following points? A. (−4, −6) B. (4, −6) C. (−4, 6) D. (2, 6)

OpenStudy (e.mccormick):

Probably the simplest way to do this is put it into y=mx+b form then try points. So start by solving for y in terms of x, AKA: get y alone. \(y-6=2(x+2)\Rightarrow y-6=2x+8\Rightarrow y=2x+14\) Now put in those x values you were given and see if you get the right y.

OpenStudy (littlebirdy):

so whats the ansers

OpenStudy (e.mccormick):

You need to do some of the work to learn. You now have the same equation in a different form: \[y=2x+14\] So take an x, put it in and see if you get the right y. The first x is -4. So put that in there, multiply it by 2, add 14, and see if you get -6 or not. If so, that is it. If not, go to the next.

OpenStudy (littlebirdy):

just please help i dont get it

OpenStudy (e.mccormick):

If I do all the work, that is not help. That is me doing the work. Help is me showing you the path and perhaps doing some of the work, which I have done.

OpenStudy (littlebirdy):

im blocking u bye

OpenStudy (e.mccormick):

Fine.

OpenStudy (littlebirdy):

just help me